How do you simplify ((e^y)(y)-(e^y)(y^2))
Answer #1

Page no.: Date: Given that le ogly) - (03) {y22 Solution We have (ed ty, - (69) 1 y 2) = g. et - y2 et =edy (1-9) = goed (1-y